OpenStudy (anonymous):

Evaluate S5 for 400 + 200 + 100 + … and select the correct answer below. 25 775 1,125 500

OpenStudy (freckles):

the series is geometric that is term/previous term equals 1/2 200/400=100/200=1/2 so 400 divided by 2 is 200 200 divided by 2 is 100 100 divided by 2 is 50 50 divided by 2 is 25 so you have your 5 terms... 400+200+100+50+25

OpenStudy (freckles):

you can also use the following formula: \[S_n=a \cdot \frac{1-r^n}{1-r} \\ \text{ where } n \text{ is the number of the terms } \\ a \text{ is the first number in the sequence } \\ r \text{ is the common ratio }\]

OpenStudy (anonymous):

thank you!

OpenStudy (freckles):

np